Is it true that if f:R→S is a ring homomorphism between two rings with unity R and S then f+1; defined by (f+1) (r) =f(r) +1 is also a ring homeomorphism from R to S?
1
Expert's answer
2020-11-26T19:42:40-0500

Let "f:R\u2192S" is a ring homomorphism between two rings with unity "R" and "S" . Consider the map"f+1: R\to S, (f+1) (r) =f(r) +1."


Since "f" is a homomorphism, "(f+1)(r+s)=f(r+s)+1=f(r)+f(s)+1". Since "S" is Abelian group according to addition operation, we conclude that

"(f+1)(r)+(f+1)(s)=f(r)+1+f(s)+1=f(r)+f(s)+1+1". If "(f+1)(r+s)=(f+1)(r)+(f+1)(s)", then "f(r)+f(s)+1=f(r)+f(s)+1+1". Using left cancelation law in Abelian group of a ring, we obtain that "1=0", and then "s=s\cdot 1=s\cdot 0=0" for any "s\in S". Therefore, if "f+1: R\to S" is a ring homomorphism, then "S=\{0\}" is a trivial ring. If "S" is a non-trivial ring, then "f+1: R\to S" is not a homomorphism.


Answer: false